Solve for x where x is a real number
log x + log (x-3) = 1

Found 2 solutions by HyperBrain, ikleyn:
Answer by HyperBrain(694)

Note: log (ab) = log a + log b.
So,
log x + log (x-3) = 1
log (x(x-3))=1
Another: 1=log 10
log (x(x-3))=log 10
taking the antilogarithms,
x(x-3)=10
Since 10=2*(5-3)=(-2-3)*-5.
Then x=5 or -2
Power up,
HyperBrain!

Answer by ikleyn(53937)

.
Solve for x where x is a real number
log x + log (x-3) = 1
~~~~~~~~~~~~~~~~~


In his post, @HyperBrain derives two solutions x = 5 and/or x = -2.

But the negative solution can not be accepted: it must be rejected,
since logarithm does not tolerate negative arguments !